5/22-23: MZ release from PSL
5/20-24: Exchanging cryo valves.
5/27-6/2: Y arm mass lock. (not done yet)
5/29: RT PC maintenance -> go to safe mode for all VIS connected Dolphin (specially BS)
6/3-6/8: Installation of TMS for ETMY
- 6/8: Engineering run
5/22AM~: tuned off cryostat cooler, finesse measurement for Xarm
5/22AM~: tuned off shield cooler, finesse measurement for Xarm
5/2: Cooling start
5/23AM~: tuned off cryostat cooler, finesse measurement for Yarm
5/17~: tuned off ETMY cryostat cooler
- Long term:
- 4/22-5/31: Yarm commissioning
- 5/27-6/30: DRMI commissioning
- 7/1-8/9: DRFPMI commissioning
- 8/12-10/4: DRFPMI ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
- Please reduce the number or speed of 16kHz channels.
- Summary meeting will be held on 6/11 11AM.
Report of work
- COM: Alignment on REFL table for camera, PD, QPD was checked.
- COM: PRMI was tried to measure OLTF.
- It was noticed that whitening filter was engaged but no anti-whitening filter was not implemented.
- LAS: NPRO will be pieced up
- CAL: Coil TF will be measured at ETMX